I need you to solve each system by using substitution.
6. y= -3x - 6
y= x-4
7. y= 5x + 8
y= -10x + 3
8.y= -4x - 8
y= 1/4x + 4
9.y= 6x - 4
y= -2x + 28
10. y= 2/5x - 3
y= 2x - 18
11. y= x - 2
2x + 2y =4
12.y= 3x - 27
y= 4x + 10y = 120
13.y= 3x - 6y =30
-y= -6x +34

Found 2 solutions by checkley71, nadinadan:
Answer by checkley71(8403)

Y=-3X-6
Y=X-4
SUBSTITUTE (X-4) FOR Y IN THE FIRST EQUATION
X-4=-3X-6
4X+3X=-6+4
7X=-2
X=-2/7 ANSWER.
Y=-2/7-4
Y=(-2-4*7)/7
Y=(-2-28)/7
Y=-30/7 ANSWER.
PROOF
-30/7=-3*-2/7-6
-30/7=6/7-6
-30/7=(6-36)/7
-30/7=-30/7

Answer by nadinadan(32)

All these equations can be solved the same way:
All you have to do is to substitute y from the first equation into the second equation:
y=-3x-6
y= x-4
y= -3x - 6 so instead of y we can write x-4=-3x-6 so we put the x on the same side, changing the sign, x+3x=4-6, so we have 4x=-2 so x=-2/4 so x= -1/2
Now in one of the equations, replace x with the value -1/2 and find y.
y=-1/2-4 so in y=x-4, y=(-1-8)/2=-9/2, so y=-9/2.
The same way you do all the exercises. Replace one variable from the first equation into the second equation and find the x and y values.